# TweedieMix: Improving Multi-Concept Fusion for Diffusion-based Image/Video Generation

### v1

@@ Abstract
Despite significant advancements in customizing text-to-image and video generation models, generating images and videos that effectively integrate multiple personalized concepts remains a challenging task. To address this, we present TweedieMix, a novel method for composing customized diffusion models during the inference phase. By analyzing the properties of reverse diffusion sampling, our approach divides the sampling process into two stages. During the initial steps, we apply a multiple object-aware sampling technique to ensure the inclusion of the desired target objects. In the later steps, we blend the appearances of the custom concepts in the de-noised image space using Tweedie’s formula. Our results demonstrate that TweedieMix can generate multiple personalized concepts with higher fidelity than existing methods. Moreover, our framework can be effortlessly extended to image-to-video diffusion models, enabling the generation of videos that feature multiple personalized concepts. Results and source code are in our anonymous project page. [^1] https://github.com/KwonGihyun/TweedieMix
@@ 2 Related Work
Building upon these advancements, there has been a growing interest in customizing T2I models using user-provided images or visual concepts. The pioneering work of Textual Inversion (Gal et al., 2022) focused on optimizing textual embeddings to represent custom concepts, enabling the generation of images that reflect these custom concepts. Subsequent studies have enhanced performance by developing extended textual embeddings (Voynov et al., 2023; Li et al., 2023) and fine-tuning model parameters (Kumari et al., 2023; Ruiz et al., 2023; Tewel et al., 2023), leading to more efficient and flexible customization options.
@@ 4 Methods
Method CLIP score User Study
@@
Textual Inversion 0.3385 0.7433 3.01 2.05 2.16
Custom Diffusion 0.3477 0.7693 2.82 2.63 2.48
Perfusion 0.3094 0.7417 2.01 2.44 1.88
Mix-of-show 0.3634 0.7984 3.82 3.77 3.56
Concept Weaver 0.3804 0.8124 4.14 4.32 4.11
TweedieMix (ours) 0.3872 0.8202 4.56 4.71 4.47
@@ 5.1 Experimental Settings
Evaluation Settings. For the evaluation dataset, we utilized the dataset proposed in the prior work, drawing from various data sources for both quantitative and qualitative analyses. For the quantitative evaluation, we selected 15 distinct concepts from the Custom Concept 101 dataset (Kumari et al., 2023), organized into five unique combinations. These concepts span a wide variety of categories, including animals, humans, natural scenes, and objects. For the qualitative analysis, we expanded the concept pool by adding three animated character concepts sourced from YouTube Blender Open Movie. All the dataset contains 5 - 8 images per each concept.
For evaluation metrics, we also borrowed the protocol provided in the recent work (Kwon et al., 2024). We evaluate our method against baseline approaches by measuring Text-alignment (Text-sim) and Image-alignment (Image-sim) using CLIP scores (Radford et al., 2021). Text-alignment calculates the cosine similarity between the CLIP embedding of the generated image and the CLIP embedding of the text prompt. To better assess our model’s ability to generate multiple concepts, we use modified the standard Image-alignment metric. This adaptation computes cosine similarity between the visual embeddings of specific concept regions and the embeddings of the corresponding target concepts. We calculate these metrics across 100 unique images generated by each model. The evaluation uses five combinations of multiple concepts, with each combination containing more than three concepts. We report the average Text-alignment and Image-alignment scores across all generated images.
@@ 5.2 Multi-concept Image Generation Results
Quantitative Evaluation. Table 1 shows the comparison results of CLIP scores and human preference user study between our model and baseline models. In terms of CLIP scores, the early models demonstrated relatively lower values for both text-image alignment scores and image-image matching scores. While the scores of the latest multi-concept models showed significant improvement over the early models, our model achieved the best quantitative scores.
@@ Appendix A More Experimental Details
Regarding sampling hyperparameters, we set the reference timestep $t_{con}$ for content-aware sampling to $0.2T$, and we found that values between $0.2T$ and $0.3T$ did not significantly affect output quality. The total timestep is set to $T$=50, and the we used image resolution of 768x768. In terms of sampling time, it takes approximately 30 seconds using a single NVIDIA RTX 3090 GPU. For resampling, we used $P=10$, and found no significant quality difference when using between 5 and 10 resampling steps.
@@
Evaluation Details. To calculate the image-alignment score, we followed the procedure outlined in the previous work, ConceptWeaver (Kwon et al., 2024). Since our generated images include multiple concepts, we couldn’t rely on whole image similarity scores. Instead, we used a text-guided segmentation model to extract concept-specific images. For example, when evaluating images with ‘[c1] dog’ and ‘[c2] cat,’ we applied the segmentation model using the prompts ‘dog’ and ‘cat’ to obtain segmented masks. We then cropped the regions containing these masks and calculated the cosine similarity between the image embeddings of the extracted sections and the corresponding concept (training) images. Since baseline methods often fail to generate all concepts, we excluded scores for images lacking all foreground objects to ensure a fair comparison.
@@
Appendix B Limitations
@@
Appendix C Additional Results
In order to further show the detailed results, we present more results in Figure 10,11. The results show that our method can generate detailed multiple custom concepts while maintaining the semantic of given text prompts.

### v2

@@ Abstract
Despite significant advancements in customizing text-to-image and video generation models, generating images and videos that effectively integrate multiple personalized concepts remains challenging. To address this, we present TweedieMix, a novel method for composing customized diffusion models during the inference phase. By analyzing the properties of reverse diffusion sampling, our approach divides the sampling process into two stages. During the initial steps, we apply a multiple object-aware sampling technique to ensure the inclusion of the desired target objects. In the later steps, we blend the appearances of the custom concepts in the de-noised image space using Tweedie’s formula. Our results demonstrate that TweedieMix can generate multiple personalized concepts with higher fidelity than existing methods. Moreover, our framework can be effortlessly extended to image-to-video diffusion models by extending the residual layer’s features across frames, enabling the generation of videos that feature multiple personalized concepts. Results and source code are in our project page. [^1] https://github.com/KwonGihyun/TweedieMix
@@ 2 Related Work
Building upon these advancements, there has been a growing interest in customizing T2I models using user-provided images or visual concepts. The pioneering work of Textual Inversion (Gal et al., 2022) focused on optimizing textual embeddings to represent custom concepts, enabling the generation of images that reflect these custom concepts. Subsequent studies have enhanced performance by developing extended textual embeddings (Voynov et al., 2023; Li et al., 2023) and fine-tuning model parameters (Kumari et al., 2023; Ruiz et al., 2023; Tewel et al., 2023), leading to more efficient and flexible customization options. Recently, framework of mutli-concept extraction (Hao et al., 2024) from single image has been proposed as an improved framework.
@@ 4 Methods
Method CLIP score DINO score User Study
@@
Textual Inversion 0.3472 0.7692 0.4420 3.01 2.05 2.16
Custom Diffusion 0.3343 0.7456 0.4965 2.82 2.63 2.48
Perfusion 0.3222 0.7182 0.4201 2.01 2.44 1.88
Mix-of-show 0.3581 0.7839 0.5211 3.82 3.77 3.56
Concept Weaver 0.3707 0.8095 0.5352 4.14 4.32 4.11
ß TweedieMix (ours) 0.3816 0.8311 0.5950 4.56 4.71 4.47
@@ 5.1 Experimental Settings
Evaluation Settings. For the evaluation dataset, we utilized the dataset proposed in the prior work, drawing from various data sources for both quantitative and qualitative analyses. For the quantitative evaluation, we selected 32 distinct concepts from the Custom Concept 101 dataset (Kumari et al., 2023), organized into 10 unique combinations. These concepts span a wide variety of categories, including animals, humans, natural scenes, and objects. For the qualitative analysis, we expanded the concept pool by adding three animated character concepts sourced from YouTube Blender Open Movie. All the dataset contains 5 - 8 images per each concept.
For evaluation metrics, we also borrowed the protocol provided in the recent work (Kwon et al., 2024). We evaluate our method against baseline approaches by measuring Text-alignment (Text-sim) and Image-alignment (Image-sim) using CLIP scores (Radford et al., 2021). Text-alignment calculates the cosine similarity between the CLIP embedding of the generated image and the CLIP embedding of the text prompt. To better assess our model’s ability to generate multiple concepts, we use modified the standard Image-alignment metric. This adaptation computes cosine similarity between the visual embeddings of specific concept regions and the embeddings of the corresponding target concepts. We calculate these metrics across 100 unique images generated by each model. The evaluation uses ten combinations of multiple concepts, with each combination containing more than three concepts. For more thorough evaluation of our concept generation performance, we calculated same image similarity score on pre-trained DINOv2 (Oquab et al., 2024) model. We report the average Text-alignment and Image-alignment scores across all generated images.
@@ 5.2 Multi-concept Image Generation Results
Quantitative Evaluation. Table 1 shows the comparison results of CLIP, DINO scores and human preference user study between our model and baseline models. In terms of CLIP and DINO scores, the early models demonstrated relatively lower values for both text-image alignment scores and image-image matching scores. While the scores of the latest multi-concept models showed significant improvement over the early models, our model achieved the best quantitative scores.
@@ 6 Conclusion
Acknowledgements This work was supported by the National Research Foundation of Korea under Grant RS-2024-00336454 and by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-11190075, Artificial Intelligence Graduate School Program, KAIST).
@@ Appendix A More Experimental Details
Regarding sampling hyperparameters, we set the reference timestep $t_{con}$ for content-aware sampling to $0.8T$, and we found that values between $0.8T$ and $0.7T$ did not significantly affect output quality. The total timestep is set to $T$=50, and the we used image resolution of 768x768. In terms of sampling time, it takes approximately 30 seconds using a single NVIDIA RTX 3090 GPU. For resampling, we used $P=10$, and found no significant quality difference when using between 5 and 10 resampling steps.
@@
Evaluation Details. To calculate the image-alignment score, we followed the procedure outlined in the previous work, ConceptWeaver (Kwon et al., 2024). Since our generated images include multiple concepts, we couldn’t rely on whole image similarity scores. Instead, we used a text-guided segmentation model to extract concept-specific images. For example, when evaluating images with ‘[c1] dog’ and ‘[c2] cat,’ we applied the segmentation model using the prompts ‘dog’ and ‘cat’ to obtain segmented masks. We then cropped the regions containing these masks and calculated the cosine similarity between the image embeddings of the extracted sections and the corresponding concept (training) images. Since baseline methods often fail to generate all concepts, we excluded scores for images lacking all foreground objects to ensure a fair comparison. For ablation study, we choose 15 different concepts from customconcept101 dataset and experimented with unique 5 different combinations. For CLIP score calculation, we also followed same protocol of quantitative evaluation. In DINO score calculation, we used the pre-trained DINO-v2-base model and reported cosine similarity between the embedding of target concept images and extracted generated object images.
@@
Figure 9: Results with various hyperparameter settings. With changing the content-aware sampling timestep, we can control how much each object contains the appearance of target custom concepts. With different number of resampling steps, we can observe the different ability of multi-object generation.
Appendix B Effect of Hyperparameters
In order to show the effect of hyperparameter change, we additionally conducted experiment by varying the key hyperparameter components: timestep for content-aware sampling $t_{con}$ and the number of resampling steps $P$. In Figure 9, we show the generated results with different settings. With varying the content-aware sampling, we can obtain more custom-concept aware outputs with using larger timestep of $t_{con}$, and the output shows more generic concepts with using smaller $t_{con}$. The result means that if we use smaller steps for custom-concept fusion, the output becomes far from target custom concepts.
With changing the resampling steps, we can observe that there is significant concept missing problem when we do not use any resampling. With using small resampling steps, we can alleviate the concept missing partially, but still there exists some of the artifacts in the generated objects. With using more resampling steps, the multi-concept outputs are more clearly generated compared to other settings.
Appendix C Time Efficiency Comparison
To compare the time consumption between our proposed method and baselines, we calculated times taken for inference on single 768x768 image. For our method, single sampling path takes 28 seconds on single RTX3090 GPU, which includes time for segmentation. For our baseline Concept Weaver, the method takes about 65 seconds, which include template generation, inversion, segmentation and sampling. For our early baseline of Mix-of-show, the method takes about 23 seconds. Although the method requires additional weight merging stage for each concept combination, we did not considered it for fair comparison. For earlier baseline of Perfusion, Custom Diffusion, Textual Inversion takes about 10 seconds for sampling, which is almost same as vanilla sampling case. Considering the superior quality of our shown results, our proposed TweedieMix shows the best efficiency on multi-concept generation.
Figure 10: Results with various style composition. Our proposed method successfully combines multiple objects with different styles, while vanilla sampling method suffer from mixed styles between the objects.
Appendix D Results on Style Composition
To verify that our proposed framework is not limited on specific custom concepts, we experimented with showing composition on various custom styles. In Figure 10, we show the output of various objects with different styles. When we see the output from vanilla sampling case, the generated outputs show difficulty in discriminating the style of each components. In our proposed method, the generated images contain separated objects with accurately reflecting the indicated style representations. The results shows that our method is not limited on specific custom concepts, but it can be extended to wider domain of image styles.
Appendix E Additional Results
Figure 11: Results on more concepts. Our model can generate high-quality multi-concept generation results when using 4 custom concepts, while our baseline Concept weaver fails to reflect the appearance of target concepts.
Figure 12: Results from Stable Video Diffusion. Our video extension can be easily adapted to other Image-to-video framework.
Results on more concepts. To show the generation performance when using more than 3 concepts, we show the comparison results using 4 different concepts in Fig. 11. The results show that our proposed method can generate realistic multi-concept aware outputs even with fusing 4 different concepts. Our baseline of Concept Weaver also can generate multiple objects, but the generated concepts often deviate from target concepts, and also suffer from severe artifacts.
Detailed comparison with baseline. In order to thoroughly compare the generation quality of our proposed method and baseline of Concept Weaver, we show the multiple generated images in Figs. 14,15,16. The results from our method successfully reflect the appearance of multiple custom concepts. In our baseline Concept Weaver, although the images contain the multiple objects, most of the generated custom concepts suffer from severe artifacts or unwanted modifications, and the generated outputs are unrealistic.
Extension to Stable Video Diffusion. To verify the flexibility of our video extension method, we applied same feature injection method on Stable Video Diffusion I2V in Figure 12. Our method can still prevent the excessive modification of target objects and generate high-quality multi-concept video.
In order to further show the detailed results, we present more results in Figure 17. The results show that our method can generate detailed multiple custom concepts while maintaining the semantic of given text prompts.
Appendix F Limitations
@@
Figure 14: Additional Comparison Results. Our model can generate high-quality multi-concept generation results on image domains.
Figure 15: Additional Comparison Results. Our model can generate high-quality multi-concept generation results compared to Concept Weaver.
Figure 16: Additional Comparison Results. Our model can generate high-quality multi-concept generation results compared to Concept Weaver.